OpenStudy (anonymous):

Nuts selling for $3.00 per pound are to be mixed with nuts seeling for $7.00 per pound. How many pounds of the more expensive nuts are needed to produce a 28-pound mixture that sells for $4.00 per pound?

OpenStudy (rulnick):

7 pounds.

OpenStudy (rulnick):

a+b=28 3a+7b=4(28) these are the equations you can use to set it up. the answer is a=21, b=7.